# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ Abstract
The increasing volume of data streams poses significant computational challenges for detecting changepoints online. Likelihood-based methods are effective, but a naive sequential implementation becomes impractical online due to high computational costs. We develop an online algorithm that exactly calculates the likelihood ratio test for a single changepoint in $p$-dimensional data streams by leveraging fascinating connections with computational geometry. This connection straightforwardly allows us to recover sparse likelihood ratio statistics exactly: that is assuming only a subset of the dimensions are changing. Our algorithm is straightforward, fast, and apparently quasi-linear. A dyadic variant of our algorithm is provably quasi-linear, being $\mathcal{O}(n\log(n)^{p+1})$ for $n$ data points and $p$ less than $3$, but slower in practice. These algorithms are computationally impractical when $p$ is larger than $5$, and we provide an approximate algorithm suitable for such $p$ which is $\mathcal{O}(np\log(n)^{\tilde{p}+1}),$ for some user-specified $\tilde{p}\leq 5.$ We derive some statistical guarantees for the proposed procedures in the Gaussian case, and confirm the good computational and statistical performance, and usefulness, of the algorithms on both empirical data and on NBA data.
Keywords computational geometry, convex hull, dynamic programming, functional pruning, online changepoint detection.
@@ 1 Introduction
In recent years, many methods have been proposed for detecting one or multiple changepoints offline or online in data streams. The reason for such a keen interest in changepoint detection methods lies in its importance for various real-world applications, including bioinformatics [1, 2], econometrics [3, 4], medicine [5, 6, 7], climate and oceanography [8, 9, 10, 11, 12], finance [13, 14], autonomous driving [15], entertainment [16, 17, 18], computer vision [19] or neuroscience [20]. Whether this is online (i.e. observations are processed as they become available) or offline (i.e. observations are processed only once the whole data is available) the ever-increasing size of the data stream raises computational challenges. Ideally, a changepoint procedure should be linear or quasi-linear in the size of the data and, for online applications, should also be able to process new observations sequentially, and as quickly as they arrive. A common idea for both offline and online parametric (multiple) changepoint model estimation, is to optimise a likelihood function or calculate likelihood ratio statistics [21, 22]. From a computational perspective this optimisation can either be done (1) heuristically using, for example, Page CUSUM [23] in an online context or Binary Segmentation and its variants [24, 14] in an offline context, or (2) exactly using dynamic programming techniques as in the Segment Neighbourhood, or Optimal Partitioning algorithms [22, 25]. Importantly, these later algorithms used with an appropriate penalty enjoy good statistical properties [26, 27, 28, 29] and have shown good properties in numerous applications [30, 31].
@@
Pruning ideas have emerged in the literature to speed up these dynamic programming algorithms without sacrificing their exactness [10, 32, 33, 34]. Two types of rules have been proposed. Inequality pruning rules, as in the PELT algorithm [10], apply to a wide range of changepoint models but are efficient only if the number of true changepoints increases quickly with the size of the data. Functional pruning, as in the pDPA, FPOP or FOCuS [32, 33, 34] algorithms, applies mostly to univariate models but is efficient even when the number of true changepoints is small in comparison to the size of the data. Interestingly, these algorithms are sequential in nature and empirically quasi-linear. This quasi-linearity was recently proven for the online detection of a single changepoint [34]. The pDPA, FPOP, and FOCuS algorithms rely on functionalisation of the problem, that is they consider the likelihood as a function of the last segment parameter. A direct extension of these ideas to a higher dimension ($p>1$) has proven difficult because it involves computing the intersection and subtraction of many convex sets in $\mathbb{R}^{p}$ [35, 36].
@@
In this paper, we study online models with one changepoint, where the likelihood can be algebraically written as a log-likelihood from the $p$-variate natural exponential family. This includes, for example, the Gaussian, Poisson, categorical and Pareto type-I distribution (assuming the minimum value is known) and also (see Section 5) non-parametric e-detectors [37]. We extend the functional pruning approach to such $p$-variate models using a linearisation argument embedding all likelihoods into a bigger space of dimension $p+2$. This idea allows us to demonstrate a link between the functional problem and problems studied in computational geometry.
@@
In particular, we show that the solutions of this extension correspond to identifying the supporting hyperplanes of a particular half-space intersection problem and for models with one changepoint demonstrate that this problem further simplifies and is dual to a lower dimensional convex hull problem of dimension $p+1$. In this later case, we further prove, assuming i.i.d data, that the number of points on this hull at time $n$ is asymptotically $\mathcal{O}(\log^{p}(n))$. Leveraging this leads to an empirically log-linear per iteration algorithm for calculating the likelihood ratio test statistics for a change at $t.$
@@
At any time-step of this algorithm, we can exactly compute the maximum likelihood ratio-statistics and maximum sparse likelihood ratio-statistics allowing only $s$ coordinates to change with $s\leq p$. In Supplementary Material H, we derive non-asymptotic bounds of these statistics to control the false alarm rate and derive a bound on the expected detection delay. While the computational cost is quasi-linear, our exact algorithms become impractical for large $p$, due to the $\log^{p}(n)$ scaling. For such cases we provide an approximation storing only an order of $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidates on the hull, where $\tilde{p}\leq p$ is chosen by the user.
The paper has the following structure. Section 2 describes the maximum likelihood framework we are using. Section 3 presents the mathematical formulation of the functional pruning problem and how it is related to a half-space intersection problem and a convex hull problem. In Section 4, the theory from Section 3 is applied to the online detection of a single changepoint with known or unknown pre-changepoint parameter. The proposed method, called MdFOCuS, is introduced in Subsection 4.3 followed with dyadic, sparse, and approximation variants in Subsection 4.4. We provide, in the Gaussian setting with known pre-change parameter, statistical guarantees (on the false alarm rate, and the detection delay) in Subsection 4.5. In Section 4.6, empirical efficiency comparisons of MdFOCuS in dimension $p$ with $p\times$FOCuS [34] and ocd [38] are performed on simulated data. In Section 5 we apply our methodology to track the performances of an NBA team and discuss how our theory could accelerate the recently proposed non-parametric e-detectors [37].
@@ 2.1 Problem set-up
We consider observations $\{y_{t}\}_{t=1,2,\dots}$ in $\mathbb{R}^{p^{\prime}}$ and wish to detect the possible presence of a changepoint and estimate its location. If there is a change, we denote the time of the change as $\tau$. We will consider likelihood-based methods for doing this, which involve postulating a model for the data within each segment between consecutive data points. Our model will be that the data point $y_{t}$ is independent and identically distributed from a distribution with density $f_{Y}(y_{t}|\theta_{1})$ before the change and $f_{Y}(y_{t}|\theta_{2})$ after the change.
At time $n$ twice the log-likelihood for the time series $\{y_{t}\}_{t=1,\dots,n}$, in terms of the pre-changepoint parameter, $\theta_{1}$ in $\mathbb{R}^{p^{\prime\prime}}$, the post-changepoint parameter, $\theta_{2}$ in $\mathbb{R}^{p^{\prime\prime}}$, and the location of a change, $\tau$, is
@@
Depending on the application one may assume that the pre-changepoint parameter $\theta_{1}$ is either known or unknown. In either case, we can get the log-likelihood for a changepoint at $\tau$ by maximising this equation over the unknown parameter, so $\theta_{2}$ if the pre-changepoint parameter is known or both $\theta_{1}$ and $\theta_{2}$ otherwise. In practice we do not know the changepoint location, so we would maximise this log-likelihood over $\tau$. We can define the maximum likelihood estimator for $\tau$, for the pre-changepoint parameter being either known or unknown respectively, as
@@
We then obtain a Generalized Likelihood Ratio (GLR) statistic [21] for a change $\tau$ by comparing with twice the log-likelihood for a model with no change. That is $2\sum_{t=1}^{n}f_{Y}(y_{t}|\theta_{1})$ if $\theta_{1}$ is known and $\max_{\theta_{1}}2\sum_{t=1}^{n}f_{Y}(y_{t}|\theta_{1})$ if it is not. Often it is appropriate (see for example [38]) to consider a sparse version of the likelihood that allows only $s$ coordinates to jump that is $s=||\theta_{1}-\theta_{2}||_{0}$, where $||\mathbf{v}||_{0}$ is equal to the number of non-zero entries of $\mathbf{v}$.
@@
$\hat{\tau}^{s}(.)=$ $\arg\max_{\tau}\left(\max_{\begin{subarray}{c}\theta_{1},\theta_{2}\\ ||\theta_{1}-\theta_{2}||_{0}=s\end{subarray}}\ell_{\tau,n}(\theta_{1},\theta_{2})\right).$ (3)
It should be clear that $\hat{\tau}^{p}(\theta_{1})=\hat{\tau}(\theta_{1})$ and $\hat{\tau}^{p}(.)=\hat{\tau}(.).$
@@
Offline, optimising Equation (1) is simple. Indeed assuming we have access to all the data it suffices to compute, for every possible changepoint location from $1$ to $n-1$, the likelihood (which can be done efficiently using summary statistics) and then pick the best value. Online, the problem is substantially more difficult as one needs to recompute the maximum for every new observation. Applying the previous offline strategy for every new observation is inefficient as it takes order $n$ to process the $n$th observation. It is infeasible for an online application, where $n$ grows unbounded, and for this reason various heuristic strategies have been proposed (see [39] for a recent review). From a computational perspective, we roughly identify two main types of heuristics. The first idea is to restrict the set of possible $(\theta_{1},\theta_{2})$ pairs to be discrete. The simplest case is to consider just one pair of $(\theta_{1},\theta_{2})$ as in the Page CUSUM strategy [23] but one typically considers a grid of values [38]. The second idea is to restrict the set of changepoint locations one considers. The simplest case is arguably to consider a grid of window lengths $w_{1},\ldots,w_{M}$ and at time $n$ compute the GLR statistic only for changes at $\tau=n-w_{1},\ldots,n-w_{M}$ [40, 41]. But this naturally introduces a trade-off between computational efficiency and statistical power: the larger the grid of windows, the larger the computational overhead [39].
@@
Recently, the FOCuS algorithm has been introduced that calculates the maximum likelihood and likelihood ratio test exactly but with a per-iteration cost that increases only logarithmically with the amount data [34, 42] - however this algorithm is only applicable for detecting changes in a univariate feature of a data stream. A key insight from the FOCuS algorithm is that to maximise the likelihood one only needs to consider changes that correspond to points on a particular convex hull in the plane. This link to geometrical features of the data is one that we leverage later in developing algorithms for detecting changes in multivariate features.
2.2 Detection of a single change in the parameter of distribution in the exponential family
@@
We briefly introduce our basic notations. We use $\mathbf{1}_{p}$ to denote a vector of length $p$ in which all elements are set to one. For any vector $\mu=(\mu^{1},\dots,\mu^{p})$ in $\mathbb{R}^{p}$, we define its $q$-norm as $||\mu||_{q}:=\left(\sum_{k=1}^{p}|\mu^{k}|^{q}\right)^{1/q}$, where $(q>1)$ in $\mathbb{R}$. For two vectors, $\mu_{1}=(\mu^{1}_{1},\dots,\mu^{p}_{1})$ and $\mu_{2}=(\mu^{1}_{2},\dots,\mu^{p}_{2})$ in $\mathbb{R}^{p}$, we define their scalar product as $\langle\mu_{1},\mu_{2}\rangle=\sum_{k=1}^{p}\mu_{1}^{k}\mu_{2}^{k}$. The symbol ||$\cdot$|| ⋅ | signifies the cardinality of $\cdot$ depending on the context.
@@
$f_{Y}(y_{t}|\theta_{k})=\exp\left\{-\frac{1}{2}\left[A^{\prime}(r(\theta_{k}))-2\langle r(\theta_{k}),s(y_{t})\rangle+B^{\prime}(s(y_{t}))\right]\right\}\,,$ (4)
@@
$f_{X}(x_{t}|\eta_{k})=\exp\left\{-\frac{1}{2}\left[A^{\prime}(\eta_{k})-2\langle\eta_{k},x_{t}\rangle+B^{\prime}(x_{t})\right]\right\}\,.$ (5)
This class encompasses many common exponential family models, see Table 1 for examples of several models of changes in the mean of $p$-variate data (where $p=p^{\prime}=p^{\prime\prime}$) and the corresponding functions $s$, $r$, $A^{\prime}$ and $B^{\prime}$. Later we will model a change in the mean ($\theta_{1}$) and variance ($\theta_{2}$) (so $p^{\prime\prime}=2$) of a Gaussian signal in $\mathbb{R}$ (so $p^{\prime}=1$). In this slightly more complex case we have $p^{\prime}\neq p=2$ and we can define $s$, $r$, $A^{\prime}$ and $B^{\prime}$ as follows
$\theta=(\theta_{1},\theta_{2}),\quad\eta=\left(\frac{\theta_{1}}{\theta_{2}},-\frac{1}{2\theta_{2}}\right),\quad s(y)=(y,y^{2}),\quad A^{\prime}(\eta)=\frac{\theta_{1}^{2}}{\theta_{2}}-\log(\theta_{2}),\quad B^{\prime}(x)=\log(2\pi).$
Let us denote the pre-changepoint natural parameter as $\eta_{1}$ and the post-changepoint natural parameter by $\eta_{2}$. We can then define the log-likelihood of our (transformed) data $x_{1},\ldots,x_{t}$, under a model with a change at $\tau$ as
@@
For any fixed $\eta_{1}$ and any $\eta_{2}$ the addition of $\sum_{t=1}^{n}\left[A^{\prime}(\eta_{2})-2\langle x_{t},\eta_{2}\rangle+B^{\prime}(x_{t})\right]$ to (6) does not change the arg-maximiser: $\arg\max_{\tau}\ell_{\tau,n}(\eta_{1},\eta_{2})$. Moreover, this addition eliminates the dependence on $n$. Hence when searching for the maximum likelihood change we can redefine $\ell_{\tau,n}(\eta_{1},\eta_{2})$ as follows:
@@
and be sure that the optimal change optimises $\ell_{\tau}(\eta_{1},\eta_{2})$ at least for a pair $(\eta_{1},\eta_{2}).$
Remark 1.
In some changepoint problems it might make sense to regularise the parameters. This can be done by adding a function of the parameter $\Omega(\eta_{1},\eta_{2})$ to the likelihood in Equation (6). For example, one could consider a lasso penalty on $\eta_{1}$ (that is, $\Omega(\eta_{1},\eta_{2})=\lambda||\eta_{1}||_{1}$ for some $\lambda>0$) to favour many $\eta_{1}$ coordinates being exactly zero. We will not study this further in the rest of the paper, but by the previous addition trick any such regularisation term could be subtracted, and therefore the rest of our argument would hold.
For all subsequent proofs, it will suffice to study the special case where $\eta_{1}$ is known. Our results for the unknown case (on $\hat{\tau}(.)$ in Theorem 5) are derived as a by-product: considering all possible $\eta_{1}$ simultaneously and realising they all lead to the same geometrical problem. Considering $\eta_{1}$ to be fixed we thus introduce the following definitions:
$\mu=\eta_{2}-\eta_{1},\quad A(\mu)=A^{\prime}(\mu+\eta_{1})-A^{\prime}(\eta_{1}),$
and rewrite the function $\ell_{\tau}(\eta_{1},\eta_{2})$ as a function of $\mu$:
@@
Distribution $\eta:=r(\theta)$ $x:=s(y)$ $A^{\prime}(\eta)$ $B^{\prime}(x)$
Gaussian (changes in mean) $\theta$ $y$ $||\eta||^{2}$ $||x||^{2}+\log(2\pi)^{p}$
Binomial $\log{\frac{\theta}{1-\theta}}$ $y$ $2m\langle\log(1+e^{\eta}),\mathbf{1}_{p}\rangle$ $-2\left\langle\log\begin{pmatrix}m\\ x\\ \end{pmatrix},\mathbf{1}_{p}\right\rangle$
Exponential $-\theta$ $y$ $-2\langle\log(-\eta),\mathbf{1}_{p}\rangle$ 0

### v3

@@ Abstract
The increasing volume of data streams poses significant computational challenges for detecting changepoints online. Likelihood-based methods are effective, but a naive sequential implementation becomes impractical online due to high computational costs. We develop an online algorithm that exactly calculates the likelihood ratio test for a single changepoint in $p$-dimensional data streams by leveraging a fascinating connection with computational geometry. This connection straightforwardly allows us to exactly recover sparse likelihood ratio statistics: that is assuming only a subset of the dimensions are changing. Our algorithm is straightforward, fast, and apparently quasi-linear. A dyadic variant of our algorithm is provably quasi-linear, being $\mathcal{O}(n\log(n)^{p+1})$ for $n$ data points and $p$ less than $3$, but slower in practice. These algorithms are computationally impractical when $p$ is larger than $5$, and we provide an approximate algorithm suitable for such $p$ which is $\mathcal{O}(np\log(n)^{\tilde{p}+1}),$ for some user-specified $\tilde{p}\leq 5$. We derive statistical guarantees for the proposed procedures in the Gaussian case, and confirm the good computational and statistical performance, and usefulness, of the algorithms on both empirical data and NBA data.
Keywords: computational geometry, convex hull, dynamic programming, online changepoint detection
@@ 1 Introduction
In recent years, many methods have been proposed for detecting one or multiple changepoints offline or online in data streams. The reason for such a keen interest in changepoint detection methods lies in its importance for various real-world applications, including bioinformatics Olshen et al. (2004); Picard et al. (2005), econometrics Bai and Perron (2003); Aue et al. (2006), medicine Bosc et al. (2003); Staudacher et al. (2005), climate and oceanography Ducré-Robitaille et al. (2003); Reeves et al. (2007), finance Andreou and Ghysels (2002); Fryzlewicz (2014), autonomous driving Galceran et al. (2017), entertainment Rybach et al. (2009); Radke et al. (2005), computer vision Ranganathan (2012) or neuroscience Jewell et al. (2022). Whether this is online (i.e. observations are processed as they become available) or offline (i.e. observations are processed only once the whole data is available) the ever-increasing size of the data stream raises computational challenges. Ideally, a changepoint procedure should be linear or quasi-linear in the size of the data and, for online applications, should also be able to process new observations sequentially, and as quickly as they arrive. A common idea for both offline and online parametric (multiple) changepoint model estimation, is to optimise a likelihood function or calculate likelihood ratio statistics Siegmund and Venkatraman (1995); Auger and Lawrence (1989). From a computational perspective this optimisation can either be done (1) heuristically using, for example, the Page-CUSUM statistic Page (1954) in an online context or Binary Segmentation and its variants Scott and Knott (1974); Fryzlewicz (2014) in an offline context, or (2) exactly using dynamic programming techniques as in the Segment Neighbourhood, or Optimal Partitioning algorithms Auger and Lawrence (1989); Jackson et al. (2005). Importantly, these later algorithms used with an appropriate penalty enjoy good statistical properties Yao (1984); Lavielle and Moulines (2000); Lebarbier (2005); Garreau and Arlot (2017) and have shown good properties in numerous applications Lai et al. (2005); Liehrmann et al. (2021).
@@
Significant speed-ups are possible by reducing the set of changepoint locations that need to be checked at each iteration of the dynamic programming recursion, a technique commonly known as pruning Killick et al. (2012); Rigaill (2015); Maidstone et al. (2017); Romano et al. (2023). Two types of rules have been proposed that prune changepoint locations that will never contribute to the full optimisation cost, achieving significant speed-ups without sacrificing optimality. Inequality pruning rules, as in the PELT algorithm Killick et al. (2012), apply to a wide range of changepoint models but are efficient only if the number of true changepoints increases quickly with the size of the data. Functional pruning, as in the pDPA, FPOP, or FOCuS algorithms Rigaill (2015); Maidstone et al. (2017); Romano et al. (2023), applies mostly to univariate models, but is efficient even when the number of true changepoints is small compared to the size of the data. Interestingly, these algorithms are sequential in nature and empirically quasi-linear. This quasi-linearity was recently proven for the online detection of a single changepoint Romano et al. (2023). The pDPA, FPOP, and FOCuS algorithms rely on a functionalisation of the problem, that is they consider the likelihood as a function of the last segment parameter. A direct extension of these ideas to a higher dimension ($p>1$) has proven difficult because it involves computing the intersection and subtraction of many convex sets in $\mathbb{R}^{p}$ Runge (2020); Pishchagina et al. (2023).
@@
In this paper, we study online models with one changepoint, where the likelihood can be algebraically written as a log-likelihood from the $p$-variate natural exponential family. This includes, for example, the Gaussian, Poisson, categorical, and Pareto type-I distribution (assuming the minimum value is known) and also the non-parametric e-detectors of Shin et al. (2022) (see Section 5). We extend the functional pruning approach to such $p$-variate models using a linearisation argument embedding all likelihoods into a bigger space of dimension $p+2$. This idea allows us to demonstrate a link between the functional problem and problems studied in computational geometry.
@@
In particular, we show that the solutions of this extension correspond to identifying the supporting hyperplanes of a particular half-space intersection problem. For models with one changepoint, we demonstrate that this problem further simplifies and is dual to a lower dimensional convex hull problem of dimension $p+1$. In this latter case, we further prove, assuming independent and identically distributed (i.i.d.) data, that the number of points on this hull at time $n$ is asymptotically $\mathcal{O}(\log^{p}(n))$. Leveraging this leads to an empirically log-linear per iteration algorithm for calculating the likelihood ratio test statistics for a change at $t$.
@@
At any time-step of this algorithm, we can exactly compute the maximum likelihood ratio-statistics and maximum sparse likelihood ratio-statistics allowing only $s$ coordinates to change with $s\leq p$. In Supplementary Material I, we derive non-asymptotic bounds of these statistics to control the false alarm rate and derive a bound on the expected detection delay. While the computational cost is quasi-linear, our exact algorithms become impractical for large $p$, due to the $\log^{p}(n)$ scaling. For such cases we
provide an approximation storing only an order of $\mathcal{O}(p\log^{\tilde{p}}(n))$ candidates on the hull, where $\tilde{p}\leq p$, chosen by the user, represents the dimension of a projective linear space.
The paper has the following structure. Section 2 describes our maximum likelihood framework. Section 3 presents the mathematical formulation of the functional pruning problem and how it is related to a half-space intersection problem and a convex hull problem. In Section 4 we develop and introduce our proposed method, called MdFOCuS, together with an approximation that scales to larger values of $p$. An important feature of both algorithms is that they calculate sets of putative changepoint locations. Given such a set it is straightforward to maximise the test statistic of interest, including under constraints on the number of components that change, or allowing different models for different components. We provide statistical guarantees of the resulting test in the Gaussian setting and compare MdFOCuS with competitor algorithms in terms of statistical and computational efficiency – showing its increased accuracy when the pre-change parameters are unknown. In Section 5 we apply our methodology to track the performances of an NBA team and discuss how our theory could accelerate the recently proposed non-parametric e-detectors Shin et al. (2022). All proofs, and links to code for this paper, are given in the Supplementary Material.
@@ 2.1 Problem set-up
We consider observations $\{y_{t}\}_{t=1,2,\dots}$ in $\mathbb{R}^{p^{\prime}}$ and wish to detect the possible presence of a changepoint and estimate its location. If there is a change, we denote the time of the change as $\tau$. We will consider likelihood-based approach that postulates a model for the data in the segments before and after the changepoint. Our model will be that the data point $y_{t}$ is i.i.d. from a distribution with density $f_{Y}(y_{t}|\theta_{1})$ before the change and $f_{Y}(y_{t}|\theta_{2})$ after the change.
At time $n$, twice the log-likelihood for the time series $\{y_{t}\}_{t=1,\dots,n}$, in terms of the pre-changepoint parameter, $\theta_{1}$ in $\mathbb{R}^{p^{\prime\prime}}$, the post-changepoint parameter, $\theta_{2}$ in $\mathbb{R}^{p^{\prime\prime}}$, and the location of a change, $\tau$, is
@@
Depending on the application one may assume that the pre-changepoint parameter $\theta_{1}$ is either known or unknown. In either case, we can get the log-likelihood for a changepoint at $\tau$ by maximising this equation over either the unknown parameter, $\theta_{2}$, or parameters, $\theta_{1}$ and $\theta_{2}$. In practice, we do not know the changepoint location, so we maximise this log-likelihood over $\tau$. We can define the maximum likelihood estimator for $\tau$, for the pre-changepoint parameter being either known or unknown respectively, as
@@
We then obtain a Generalized Likelihood Ratio (GLR) statistic Siegmund and Venkatraman (1995) for a change $\tau$ by comparing it with twice the log-likelihood for a model with no change. That is $2\sum_{t=1}^{n}\log f_{Y}(y_{t}|\theta_{1})$ if $\theta_{1}$ is known and $\max_{\theta_{1}}2\sum_{t=1}^{n}\log f_{Y}(y_{t}|\theta_{1})$ if it is not. Often it is appropriate (see, for example, Chen et al., 2022) to consider a sparse version of the likelihood that allows only $s$ coordinates to jump, that is adding the constraint $s=||\theta_{1}-\theta_{2}||_{0}$, where $||\mathbf{v}||_{0}$ is equal to the number of non-zero entries of $\mathbf{v}$. The estimator of $\tau$ in this case, assuming the pre-change mean is known, is
@@
with $\hat{\tau}^{s}(.)$ defined analogously for the case where the pre-change mean is unknown. Clearly, $\hat{\tau}^{p}(\theta_{1})=\hat{\tau}(\theta_{1})$ and $\hat{\tau}^{p}(.)=\hat{\tau}(.)$.
@@
Offline, optimising Equation (1) is simple. Indeed, assuming we have access to all the data, it suffices to compute, for every possible changepoint location from $1$ to $n-1$, the likelihood (which can be done efficiently using summary statistics) and then pick the best value. Online, the problem is substantially more difficult as one needs to recompute the maximum for every new observation. Applying the previous offline strategy for every new observation is inefficient as it takes a computational cost of order $n$ to process the $n$th observation. It is infeasible for an online application, where $n$ grows unbounded, and for this reason, various heuristic strategies have been proposed (see Wang and Xie, 2022, for a recent review). From a computational perspective, we roughly identify two main types of heuristics. The first idea is to restrict the set of possible $(\theta_{1},\theta_{2})$ pairs to be discrete. The simplest case is to consider just one pair of $(\theta_{1},\theta_{2})$ as in the Page CUSUM strategy Page (1954), but one typically considers a grid of values Chen et al. (2022). The second idea is to restrict the set of changepoint locations one considers. For example, to consider a grid of window lengths $w_{1},\ldots,w_{M}$ and at time $n$ compute the GLR statistic only for changes at $\tau=n-w_{1},\ldots,n-w_{M}$ Eiauer and Hackl (1978); Meier et al. (2021). But this naturally introduces a trade-off between computational efficiency and statistical power: the larger the grid of windows, the larger the computational overhead Wang and Xie (2022).
@@
Recently, the FOCuS algorithm has been introduced that calculates the GLR test exactly but with a per-iteration cost that increases only logarithmically with the amount of data Romano et al. (2023); Ward et al. (2023) - however, this algorithm is only applicable for detecting changes in a univariate feature of a data stream. A key insight from the FOCuS algorithm is that to maximise the likelihood, one only needs to consider changes that correspond to points on a particular convex hull in the plane. This link to geometrical features of the data is one that we leverage to develop algorithms for detecting changes in multivariate features.
2.2 Detection of a single change in the parameter of a distribution in the exponential family
@@
We briefly introduce some notation. We use $\mathbf{1}_{p}$ to denote a vector of length $p$ in which all elements are set to one. For any vector $\mu=(\mu^{1},\dots,\mu^{p})$ in $\mathbb{R}^{p}$, we define its $q$-norm as $||\mu||_{q}:=\left(\sum_{k=1}^{p}|\mu^{k}|^{q}\right)^{1/q}$, where $q\geq 1$ in $\mathbb{R}$. For two vectors, $\mu_{1}=(\mu^{1}_{1},\dots,\mu^{p}_{1})$ and $\mu_{2}=(\mu^{1}_{2},\dots,\mu^{p}_{2})$ in $\mathbb{R}^{p}$, we define their scalar product as $\langle\mu_{1},\mu_{2}\rangle=\sum_{k=1}^{p}\mu_{1}^{k}\mu_{2}^{k}$. The symbol ||$\cdot$|| ⋅ | signifies the cardinality of $\cdot$.
@@
$f_{Y}(y_{t}|\theta_{k})=\exp\left\{-\frac{1}{2}\left(A^{\prime}(r(\theta_{k}))-2\langle s(y_{t}),r(\theta_{k})\rangle+B^{\prime}(s(y_{t}))\right)\right\}\,,$ (3)
@@
$f_{X}(x_{t}|\eta_{k})=\exp\left\{-\frac{1}{2}\left(A^{\prime}(\eta_{k})-2\langle x_{t},\eta_{k}\rangle+B^{\prime}(x_{t})\right)\right\}\,.$ (4)
This class encompasses many common exponential family models, see Table 4 in the Supplementary Material. There, we also show how a change in mean and variance of a Gaussian signal in $\mathbb{R}$ can be expressed in this form.
We can then define the log-likelihood of our (transformed) data $x_{1},\ldots,x_{t}$, under a model with a change at $\tau$ in terms of $\eta_{1}$ and $\eta_{2}$ as
@@
For any fixed $\eta_{1}$ and any $\eta_{2}$ the addition of $\sum_{t=1}^{n}\left[A^{\prime}(\eta_{2})-2\langle x_{t},\eta_{2}\rangle+B^{\prime}(x_{t})\right]$ to (5) does not change $\arg\max_{\tau}\ell_{\tau,n}(\eta_{1},\eta_{2})$. Moreover, this addition eliminates the dependence on $n$. Hence, when searching for the maximum likelihood change, we can redefine $\ell_{\tau,n}(\eta_{1},\eta_{2})$ as follows:
@@
and be sure that the optimal change optimises $\ell_{\tau}(\eta_{1},\eta_{2})$ at least for a pair $(\eta_{1},\eta_{2})$. Sometimes it might make sense to regularise the parameters. This can be done by adding a function of the parameter $\Omega(\eta_{1},\eta_{2})$, such as $\Omega(\eta_{1},\eta_{2})=\lambda(||\eta_{1}||_{1}+||\eta_{2}||_{1})$ for some $\lambda>0$, to (5). We will not study this further, but by the previous addition trick, any such regularisation term could be subtracted, and the rest of our argument would hold.
Throughout, we will primarily focus on the case where $\eta_{1}$ is known. Our results for the case where it is unknown are derived as a by-product: considering all possible $\eta_{1}$ simultaneously and realising they all lead to the same geometrical problem. Considering $\eta_{1}$ to be fixed we thus introduce the following definitions: $\mu=\eta_{2}-\eta_{1},\quad A(\mu)=A^{\prime}(\mu+\eta_{1})-A^{\prime}(\eta_{1})$, and rewrite the function $\ell_{\tau}(\eta_{1},\eta_{2})$ as a function of $\mu$: